# FedMeld: A Model-dispersal Federated Learning Framework for Space-ground Integrated Networks

### v1

@@ IV FedMeld Optimization
In this section, the results from the preceding convergence analysis are applied to the FedMeld optimization. First, we formulate a joint optimization problem of SC-MR to minimize the model loss under FedMeld within a fixed span of training time. Then, we demonstrate that the problem can be decomposed into two consecutive subproblems, which can be optimally solved using the algorithm we proposed.
@@ IV-A Problem Formulation
where the objective (18a) is to optimize the convergence bound. Constraint (18b) ensures that the sum of idle duration and additional training time equals the satellite flight time between two adjacent areas. Constraint (18c) ensures that the total training time does not exceed the maximum constraint. Constraint (18d) limits the maximal round gap of the mixed models, which has been discussed in detail earlier in Theorem 1.
@@ IV-B Equivalent Problem Decomposition
The feasibility of $\mathcal{P}1$ indicates the right side of (18d) is large enough. It is clear that the objective function $f\left(\delta,\alpha\right)$ of $\mathcal{P}1$ is monotonically increasing with respect to $\delta$. Simultaneously, the right side of (18d) is a monotonically decreasing function with respect to $\alpha$. Thus, solving $\mathcal{P}2$, which can obtain the minimal value of $\delta$ (i.e., $\delta^{*}$), can allow for a larger range of values for $\alpha$. It can increase the likelihood of $\alpha$ reaching its optimal solution. Conditioned on $\delta^{*}$, the optimal mixing ratio policy is simplified to $\mathcal{P}3$. ∎
@@ IV-C2 Optimal Solution of $\alpha$
First, we discuss the feasible set of $\alpha$. By taking the derivative, we can find that if the monotonic decreasing function $m\left(\alpha\right)$ satisfies $m\left(\alpha\right)\geq 2$, i.e., $\alpha\leq\frac{1}{2}$, the right side of (18d) is a non-decreasing function of $t$ Therefore, as long as ${\delta^{*}}E\leq\frac{{m\left(\alpha\right){{\left[{\left({K+{\delta^{*}}}\right)E+\gamma}\right]}^{2}}}}{{{{\left[{\left({K+{\delta^{*}}}\right)E+\gamma-1}\right]}^{2}}+m\left(\alpha\right)\left[{\left({K+{\delta^{*}}}\right)E+\gamma}\right]}}$ holds, which is equivalent to
@@
then (18d) can be always satisfied. Let $\tilde{\alpha}$ be the value of $\alpha$ which makes (22) an equality. Thus, the feasible set of $\alpha$ is $\alpha\in\left[0,\min\left\{\tilde{\alpha},\frac{1}{2}\right\}\right]$.
@@
Remark 4.
From (21), we have the following observation. When the training latency requirements become more stringent, or the satellite distribution becomes denser, $\delta^{\ast}$ increases. This indicates that the difference in the global round interval of model mixing between adjacent regions becomes larger.
Remark 5.
The large non-IID degree of data $\Gamma$ leads to bigger $\zeta_{1}$. When $\zeta_{1}$ is large enough, $\alpha_{0}$ will be smaller, making $\alpha^{*}<\frac{1}{2}$ possible. This implies that when the non-IID degree of data is larger, the mixing ratio of historical models from adjacent areas should be less.
@@ V-A Experiment Settings
We conduct image classification tasks using the CIFAR-10 [30] and MNIST [31] datasets, employing the ResNet-18 architecture [32]. The CIFAR-10 dataset comprises 50,000 training samples and 10,000 testing samples across 10 categories, while MNIST contains 60,000 training samples and 10,000 testing samples with labels ranging from 0 to 9. Each client possesses an equal number of images, with distinct images allocated to each device. The computational capacity of each client is $h_{j}=15.11$ TFLOPS. In each global round, 80% of clients from each cluster are randomly selected to participate in the training process. Furthermore, we set $E=5$ [21] and $|\varsigma|=64$ [33]. The initial learning rate is determined using a grid search method [29]. The maximum tolerable delay, $T_{\max}$, is 32 hours for CIFAR-10 and 20 hours for MNIST. Additional key wireless parameters are provided in Table I [34].
TABLE I: Main parameter setting
@@
TABLE II: Comparison of typical federated learning framework with $M$ clusters and $U$ participating clients in one global round
@@
Model aggregation/exchanging method Ground station ISL N.A.
@@
  • HFL [10]. In each global round, each serving satellite activates a subset of clients within its coverage area to perform $E$ local iterations before aggregating the models from these clients. After $K+\delta$ global rounds, the satellites transmit the aggregated models to a ground station for global aggregation. Upon re-establishing contact with the ground station, satellites receive the updated global model. In our experiments, two ground stations are used for global aggregation, with parameter sharing facilitated through ground optical fibers.
@@ V-B Communication Costs
Before evaluating the learning performance, Table II provides a comparison between FedMeld and three baseline algorithms with respect to learning type, model aggregation methods, and communication costs per global round. This analysis considers a scenario with $M$ clusters and $U$ participating clients. Here, $\min$ refers to the optimal conditions for PFL and Ring Allreduce, where all satellites serving the clusters can establish ISLs with their neighbors. When these serving satellites cannot set up ISLs, the models of each area must be sent to the ground station. Therefore, $\max$ denotes the worst-case scenario where any two satellites are not visible to each other. We can conclude that the proposed FedMeld is an infrastructure-independent framework with the lowest communication costs.
@@ V-C Learning Performance
In this part, we evaluate the learning performance of various FL frameworks across two datasets under three distinct data distribution settings. Fig. 3 and Fig. 4 illustrate the learning performance of the FedMeld compared to three baseline schemes across different data distribution scenarios in each dataset.
@@
From the experiment results, the following observations can be made: 1) The proposed FedMeld consistently achieves the highest model accuracy among all frameworks, demonstrating the superior performance of our approach. Compared to HFL, this decentralized FL scheme allows adjacent regions to mix parameters directly and update models more flexibly, avoiding the excessive time required for model aggregation at ground stations. The advantages of FedMeld over PFL and Ring Allreduce in terms of model accuracy stem from two key aspects: First, the careful design of mixing ratio and global round interval between adjacent regions. Second, FedMeld avoids reliance on ISLs that are not yet fully deployed across the entire constellation. 2) Ring Allreduce exhibits the fastest convergence speed. This is because, after the same number of global rounds, Ring Allreduce completes a full global average, whereas FedMeld and P-FedAvg perform local parameter exchanges between neighboring regions. 3) The model accuracy of P-FedAvg declines markedly as the non-IID degree of the data increases. 4) The accuracy of HFL is comparable to that of FedMeld. However, HFL always has the longest training time. This is because the aggregation frequency in HFL is determined by the interval between consecutive passes of all serving satellites over the ground stations.
@@ V-D Effects of Satellite Number
Figure 5: Effect of satellite number on test accuracy and training time on CIFAR-10.
@@ VI Conclusion
In this paper, we propose the novel FedMeld framework for SGINs. By leveraging the predictable movement patterns and SCF capabilities of satellites, the infrastructure-free FedMeld enables parameter aggregation across different terrestrial regions without relying on ground stations or ISLs. Our findings indicate that the optimal global round interval of model mixing between adjacent areas is highly influenced by latency constraints and handover frequency, while the ideal mixing ratio of historical models from adjacent regions is determined by the degree of non-IID data distribution.
This work pioneers the concept of an infrastructure-free FL framework within the context of space-ground Fluid AI. The current FedMeld algorithm can be extended to more complex scenarios, such as implementing region-specific round intervals and adaptive mixing ratios. Practical challenges, such as balancing learning performance with handover and launch costs, merit further exploration. Additionally, investigating efficient inference strategies and model downloading mechanisms for SGINs represents a promising avenue for future research.
@@ Appendix A Proof of Theorem 1
$\begin{split}&\mathbb{E}{\left\|{{{\mathbf{g}}_{t}}-{{\overline{\mathbf{g}}}_{t}}}\right\|^{2}}\\ &={\left\|{\sum\limits_{i\in{\mathcal{M}}}{\sum\limits_{j\in{{\mathcal{N}}_{i}}}{\frac{1}{{M{N_{i}}}}\left[{\nabla{F_{j}}\left({{\mathbf{w}}_{t,j}^{s},\varsigma_{t,j}^{s}}\right)-\nabla{F_{j}}\left({{{\mathbf{w}}_{t,j}}}\right)}\right]}}}\right\|^{2}}\\ &\leq\sum\limits_{i\in{\mathcal{M}}}{\sum\limits_{j\in{{\mathcal{N}}_{i}}}{\frac{{\sigma_{j}^{2}}}{{{{\left({M{N_{i}}}\right)}^{2}}}}}}.\\ \end{split}$ (26)
@@ Appendix B Proof of Lemma 2
We introduce an auxiliary variable $\tilde{t}$. When $t\notin{\mathcal{I}}_{E}$, then $\tilde{t}=\left\lfloor{t/E}\right\rfloor E$. When $t\in{\mathcal{I}}_{E}$, $\tilde{t}=t-E$ holds. Then, the update rule of local model can be written as
${\mathbf{w}_{t,j}}=\left\{\begin{array}[]{l}{\mathbf{w}_{\tilde{t},j}}-\sum\limits_{s=\tilde{t}}^{t-1}{{\eta_{s}}{\mathbf{g}_{s,j}}},\;{\text{if}}\;t\notin{\mathcal{I}}_{E},\\ \frac{{1-\alpha{A_{t}}}}{{{N_{i}}}}\sum\limits_{j\in{{{\cal N}}_{i}}}{\left({{\mathbf{w}_{\tilde{t},j}}-\sum\limits_{s=\tilde{t}}^{t-1}{{\eta_{s}}{\mathbf{g}_{s,j}}}}\right)}+\alpha{A_{t}}{{\overline{\mathbf{w}}}_{t-\delta E,i-1}},\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\;{\text{if}}\;t\in{\mathcal{I}}_{E},\\ \end{array}\right.$ (34)
@@ Appendix C Proof of Theorem 2
First, we define the following variables $C_{1}-C_{5}$ for notational brevity. That is, ${C_{1}}=\left({1+b}\right){\left({1-\alpha}\right)^{2}}$, ${C_{2}}=\left({1+b}\right){\alpha^{2}}$, ${C_{3}}=4\left({1+\frac{1}{b}}\right){\left({1-\alpha}\right)^{2}}{\left({E-1}\right)^{2}}{G^{2}}$, ${C_{4}}=\left({1+b}\right)$, and ${C_{5}}=4\left({1+\frac{1}{b}}\right){\left({E-1}\right)^{2}}{G^{2}}$. When $R-1$ is the step of different area mixing models, i.e., $A_{R-1}=1$, we have the following set of inequalities from step ${R-1-\left({K+\delta}\right)E+E}$ to $R-1$ based on (12).
$\left\{\begin{gathered}{Q_{R-1}}\leq{C_{1}}{Q_{R-1-E}}+{C_{2}}{Q_{R-1-\delta E}}+{C_{3}}\eta_{R-1}^{2},\hfill\\ {Q_{R-1-E}}\leq{C_{4}}{Q_{R-1-2E}}+{C_{5}}\eta_{R-1-E}^{2},\hfill\\ \ldots\hfill\\ {Q_{R-1-\left({K+\delta}\right)E+E}}\leq{C_{4}}{Q_{R-1-\left({K+\delta}\right)E}}+{C_{5}}\eta_{R-1-\left({K+\delta}\right)E+E}^{2}.\hfill\\ \end{gathered}\right.$
@@
Assume that $R-1=q\left(K+\delta\right)E$, where $q$ denotes the total model mixing times from $t=0$ to $R-1$. With the similar form of (38), we have the following inequalities of $Q_{t}$ when $A_{t}=1$.
$\left\{\begin{gathered}{Q_{R-1}}\leq{\kappa_{1}}{Q_{R-1-\left({K+\delta}\right)E}}+{\kappa_{2}}\eta_{T-1-\left({K+\delta}\right)E+E}^{2},\hfill\\ {Q_{R-1-\left({K+\delta}\right)E}}\leq{\kappa_{1}}{Q_{R-1-2\left({K+\delta}\right)E}}+{\kappa_{2}}\eta_{R-1-2\left({K+\delta}\right)E+E}^{2},\hfill\\ ...\hfill\\ {Q_{R-1-\left({q-1}\right)\left({K+\delta}\right)E}}\leq{\kappa_{1}}{Q_{R-1-q\left({K+\delta}\right)E}}+{\kappa_{2}}\eta_{R-1-q\left({K+\delta}\right)E+E}^{2},\hfill\\ \end{gathered}\right.$

### v2

@@ IV FedMeld Optimization
In this section, the results from the preceding convergence analysis are applied to the FedMeld optimization. First, we formulate a joint optimization problem of SC-MR to minimize the upper bound of the training loss under FedMeld within a fixed span of training time. Then, we demonstrate that the problem can be decomposed into two consecutive subproblems, which can be optimally solved using the algorithm we proposed.
@@ IV-A Problem Formulation
where the objective (19a) is to optimize the convergence bound. Constraint (19b) ensures that the sum of idle duration and additional training time equals the satellite flight time between two adjacent areas. During the stage from when the SCF satellite leaves the visible range of area $i$ until completing model aggregation for area $\left(i+1\right)$, we define two parameters: $T_{i,i+1}^{{\rm{fly}}}$ and $T_{i,i+1}^{{\rm{idle}}}$. Here, $T_{i,i+1}^{{\rm{fly}}}$ is the total duration of this process, which can be determined using publicly available Two-Line Element sets (TLEs), and $T_{i,i+1}^{{\rm{idle}}}$ is the idle duration of clients in the region $(i+1)$, i.e., they are neither training with SCF satellites nor non-SCF satellites. Then, we have $T_{i,i+1}^{{\rm{idle}}}+\sum\limits_{k^{\prime}=k+1}^{k+\delta}{T_{k^{\prime},i+1}^{{\rm{total}}}}=T_{i,i+1}^{{\rm{fly}}}$ for any $k-K\bmod(K+\delta)=0$. Constraint (19c) ensures that the total training time does not exceed the maximum constraint. Constraint (19d) limits the maximum round gap of the mixed models, which has been discussed in Theorem 1.
@@ IV-B Equivalent Problem Decomposition
The feasibility of $\mathcal{P}1$ indicates the right side of (19d) is large enough. It is clear that the objective function $f\left(\delta,\alpha\right)$ of $\mathcal{P}1$ is monotonically increasing with respect to $\delta$. Simultaneously, the right side of (19d) is a monotonically decreasing function with respect to $\alpha$. Thus, solving $\mathcal{P}2$, which can obtain the minimum value of $\delta$ (i.e., $\delta^{*}$) under its constraints, leads to the largest feasible range for $\alpha$, which must contain the optimal solution of $\alpha$ (i.e., $\alpha^{*}$). Obviously, plugging $\delta^{*}$ into $\mathcal{P}3$ and then solving $\mathcal{P}3$ leads to the global optimal solution. ∎
@@ IV-C1 Optimal Solution of $\delta$
With (22), we have the following observation.
Remark 5 (Discussion of the optimal model staleness indicator).
When the training latency requirements become more stringent, or the satellite distribution becomes denser, $\delta^{\ast}$ increases. This indicates that the difference in the global round interval of model mixing between adjacent regions becomes larger.
@@ IV-C2 Optimal Solution of $\alpha$
First, we discuss the feasible set of $\alpha$. By taking the derivative, we can find that if the monotonic decreasing function $m\left(\alpha\right)$ satisfies $m\left(\alpha\right)\geq 2$, i.e., $\alpha\leq\frac{1}{2}$, the right side of (19d) is a non-decreasing function of $t$ . Therefore, as long as ${\delta^{*}}E\leq\frac{{m\left(\alpha\right){{\left[{\left({K+{\delta^{*}}}\right)E+\gamma}\right]}^{2}}}}{{{{\left[{\left({K+{\delta^{*}}}\right)E+\gamma-1}\right]}^{2}}+m\left(\alpha\right)\left[{\left({K+{\delta^{*}}}\right)E+\gamma}\right]}}$ holds, which is equivalent to
@@
then (19d) can be always satisfied. Let $\tilde{\alpha}$ be the value of $\alpha$ which makes (23) an equality, i.e., $m\left(\tilde{\alpha}\right)=\frac{{{\delta^{*}}E{{\left[{\left({K+{\delta^{*}}}\right)E+\gamma-1}\right]}^{2}}}}{{\left[{\left({K+{\delta^{*}}}\right)E+\gamma}\right]\left({KE+\gamma}\right)}}$. Thus, the feasible set of $\alpha$ is $\alpha\in\left[0,\min\left\{\tilde{\alpha},\frac{1}{2}\right\}\right]$.
@@
Remark 6 (Discussion of the optimal mixing ratio).
The large non-IID degree of data $\Gamma$ leads to bigger $\zeta_{1}$. When $\zeta_{1}$ is large enough, $\alpha_{0}$ will be smaller, making $\alpha^{*}<\frac{1}{2}$ possible. This implies that when the non-IID degree of data is higher, the mixing ratio of historical models from adjacent areas should be less. Intuitively, when regional data distributions are highly heterogeneous, the historical model parameters from another region are not only stale but also biased toward a divergent data domain. Assigning a large weight to such a model may distort the optimization trajectory of the global model, resulting in slower or unstable convergence. Therefore, a smaller $\alpha$ helps reduce the influence of stale or biased historical models, thereby enhancing the stability of convergence.
@@ V-A Experiment Settings
Similar to previous works [19, 20, 17, 44], we conduct image classification tasks using the CIFAR-10 [45] and MNIST [46] datasets, employing the ResNet-18 architecture [47]. Given that image classification is a key task in remote applications like remote sensing and healthcare, these datasets provide a relevant and practical basis for validating our proposed FedMeld. The CIFAR-10 dataset comprises 50,000 training samples and 10,000 testing samples across 10 categories, while MNIST contains 60,000 training samples and 10,000 testing samples with labels ranging from 0 to 9. Each client possesses an equal number of images, with distinct images allocated to each device. The necessary parameters for executing the optimization problem are estimated using the method in [48]. The computational capacity of each client is $h_{j}=15.11$ TFLOPS. In each global round, 80% of clients from each cluster are randomly selected to participate in the training process. Furthermore, we set $E=5$ [25] and $|\varsigma|=64$ [49]. The initial learning rate is determined using a grid search method [43]. The maximum tolerable delay, $T_{\max}$, is 32 hours for CIFAR-10 and 20 hours for MNIST. Additional key wireless parameters are provided in Table II [50].
TABLE II: Main parameter setting
@@
TABLE III: Comparison of typical federated learning framework with $M$ clusters and $U$ participating clients in one global round
@@
Extra infrastructure Ground station ISL None
@@
  • Hierarchical FL (HFL) [17]. In each global round, each serving satellite activates a subset of clients within its coverage area to perform $E$ local iterations before aggregating the models from these clients. After $K+\delta$ global rounds, the satellites transmit the aggregated models to a ground station for global aggregation. Upon re-establishing contact with the ground station, satellites receive the updated global model. In our experiments, two ground stations are used for global aggregation, with parameter sharing facilitated through ground optical fibers.
@@ V-B Communication Costs
Before evaluating the learning performance, Table III provides a comparison between FedMeld and three baseline algorithms with respect to learning type, model aggregation methods, and communication costs per global round. This analysis considers a scenario with $M$ clusters and $U$ participating clients. In addition, $l_{\rm{U2S}}$, $l_{\rm{S2G}}$, and $l_{\rm{ISL}}$ denote the numbers of user-to-satellite links, satellite-to-ground-station links, and inter-satellite links, respectively. The $\min$ and $\max$ in Table III mainly target the SGI-FL frameworks using ISLs for parameter exchange. That is, PFL and Ring AllReduce. Here, $\min$ refers to the optimal condition where the satellites serving any two adjacent clusters can directly establish ISLs, thereby minimizing the number of communication links required. In contrast, $\max$ denotes the worst-case condition where the satellites serving any two adjacent clusters cannot establish ISLs. In this case, if two satellites wish to exchange models, each must first send its model to a ground station and then receive the counterpart model via the ground station. Consequently, a model exchange that would otherwise require one ISL now requires two satellite-to-ground station link transmissions. We can conclude that the proposed FedMeld is an infrastructure-independent framework with the lowest communication costs.
@@ V-C Learning Performance
In this part, we evaluate the learning performance of various FL frameworks across two datasets under three distinct data distribution settings. Fig. 3 and Fig. 4 illustrate the learning performance of the FedMeld compared to three baseline schemes across different data distribution scenarios in each dataset. These three data distribution settings emulate different levels of task and data heterogeneity across clients and clusters.
@@
From the experiment results, the following observations can be made: 1) The proposed FedMeld consistently achieves the highest model accuracy among all frameworks, demonstrating the superior performance of our approach. Compared to HFL, this decentralized FL scheme allows adjacent regions to mix parameters directly and update models more flexibly, avoiding the excessive time required for model aggregation at ground stations. The advantages of FedMeld over PFL and Ring Allreduce in terms of model accuracy stem from two key aspects: First, the careful design of mixing ratio and global round interval between adjacent regions. Second, FedMeld avoids reliance on ISLs that are not yet fully deployed across the entire constellation. 2) Ring Allreduce exhibits the fastest convergence speed. This is because, after the same number of global rounds, Ring Allreduce completes a full global average, whereas FedMeld and P-FedAvg perform local parameter exchanges between neighboring regions. However, despite its faster convergence, Ring Allreduce introduces significantly higher communication overhead, as it requires frequent utilization of ISLs to mix models across neighboring regions. The reliance on extensive ISL communication can become a bottleneck, particularly in large-scale or bandwidth-constrained satellite networks. 3) The model accuracy of P-FedAvg declines markedly as the non-IID degree of the data increases. 4) The accuracy of HFL is comparable to that of FedMeld. However, HFL always has the longest training time. This is because the aggregation frequency in HFL is determined by the interval between consecutive passes of all serving satellites over the ground stations.
@@ V-D Effects of Satellite Number
Figure 5: Effect of number of satellites participating in training per orbit on test accuracy and training time on CIFAR-10.
@@ V-E Effects of Latency Constraint
V-F Effect of Data Heterogeneity
Fig. 9 and Fig. 10 illustrate the impact of intra-cluster and inter-cluster data heterogeneity on the test accuracy and optimal mixing ratio for CIFAR-10 and MNIST, respectively. In these experiments, the degree of data heterogeneity is controlled by the Dirichlet concentration parameters, which determine the skewness of the data distributions across clusters and among users within each cluster [42, 51]. The concentration parameters are varied in $\left\{0.5,1,2,5,10\right\}$, covering a range from highly non-IID to nearly IID distributions. A smaller concentration parameter corresponds to a more imbalanced (highly non-IID) data distribution, whereas a larger parameter leads to a more uniform (approaching IID) allocation of samples. As shown in the figures, when the heterogeneity increases, i.e., the Dirichlet concentration parameter decreases, the optimal mixing ratio consistently decreases, confirming our Remark 6 that greater distribution disparity requires weaker inter-cluster or inter-user model mixing to mitigate negative transfer across dissimilar datasets. Meanwhile, a lower degree of heterogeneity generally yields higher test accuracy, as more balanced data partitions enable stable aggregation and faster convergence. These results collectively demonstrate that the proposed adaptive mixing mechanism can dynamically adjust to different levels of data heterogeneity.
(a) IID clusters
(b) Non-IID clusters
Figure 9: Effect of data heterogeneity with different non-IID degrees on CIFAR-10.
(a) IID clusters
(b) Non-IID clusters
Figure 10: Effect of data heterogeneity with different non-IID degrees on MNIST.
@@ VI Conclusion
In this paper, we propose the novel FedMeld framework for SGINs. By leveraging the predictable movement patterns and SCF capabilities of satellites, the infrastructure-free FedMeld enables parameter aggregation across different terrestrial regions without relying on ground stations or ISLs. Our findings indicate that the optimal global round interval of model mixing between adjacent areas is highly influenced by latency constraints and handover frequency, while the optimal mixing ratio of historical models from adjacent regions is determined by the degree of non-IID data distribution.
This work pioneers the concept of an infrastructure-free FL framework within the context of SGI-FL. The current FedMeld algorithm can be extended to more complex scenarios, such as implementing region-specific round intervals and adaptive mixing ratios. Practical challenges, such as balancing learning performance with handover and launch costs, merit further exploration. Additionally, investigating efficient inference strategies and model downloading mechanisms for SGINs represents a promising avenue for future research.
@@ Appendix A Proof of Theorem 1
[…]
@@ Appendix B Proof of Lemma 2
[…]
@@ Appendix C Proof of Theorem 2
First, we define the following variables $C_{1}-C_{5}$ for notational brevity. That is, ${C_{1}}=\left({1+b}\right){\left({1-\alpha}\right)^{2}}$, ${C_{2}}=\left({1+b}\right){\alpha^{2}}$, ${C_{3}}=4\left({1+\frac{1}{b}}\right){\left({1-\alpha}\right)^{2}}{\left({E-1}\right)^{2}}{G^{2}}$, ${C_{4}}=\left({1+b}\right)$, and ${C_{5}}=4\left({1+\frac{1}{b}}\right){\left({E-1}\right)^{2}}{G^{2}}$. When R1R-1 is the step of different area mixing models, i.e., $A_{R-1}=1$, we have the following set of inequalities from step $R-1-\left({K+\delta}\right)E+E$} to R1R-1 based on (13).
[…]
@@
Assume that $R-1=q\left(K+\delta\right)E$, where $q$ denotes the total model mixing times from $t=0$ to R1R-1. With the similar form of (39), we have the following inequalities of $Q_{t}$ when $A_{t}=1$.
[…]